OpenStudy (anonymous):

Which statement is true about the difference 2radical7 - radical 28

OpenStudy (anonymous):

\[2\sqrt{7}-\sqrt{28}\] 1. it is rational and equal to -2 2. It is rational and equal to 0 3. It is irrational and equal to \[-2\sqrt{7}\] 4. It is irrational and equal to \[\sqrt{7}\]